# Digital Twin based Automatic Reconfiguration of Robotic Systems in Smart Environments

### v1

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Robotic systems have achieved significant penetration in smart environments, including smart cities, smart manufacturing, and smart agriculture. Smart unmanned vehicles that deliver services inside cities, drones that monitor critical infrastructure of the city (i.e. traffic), UAVs that assist tourists, or even robotics that help pedestrians and especially persons with mobility impairments, exemplify the integration of robotics into smart city ecosystems  [1]. Moreover, robots have become an integral part of modern agricultural infrastructure, automating tasks and assisting workers in their daily operations  [2]. All these robotic systems require precise digital control, typically provided by industrial-grade controllers similar to those used in other smart infrastructures like buildings and factories. Another characteristic of smart environments is their changeability over time, especially regarding the choreographic aspect. It is common for a city to construct new buildings, roads or parks causing changes in the landscape, while this also stands true in the agricultural sector with changes in crop fields as well as relevant processing and storage infrastructures [3]. Changes in the topology of smart environments demand quick and dependable controller reconfiguration, so that they can adapt to changes and effectively plan and execute robotic system movements.
The concept of Digital Twins (DTs) has gained significant traction in the robotics domain, offering real-time synchronization between physical assets and their virtual counterparts  [4]. Early applications focused on offline simulation and predictive maintenance in industrial robotics [5], but recent work emphasizes their role in runtime monitoring, control, and autonomy. For instance, Lu et al. [6] proposed a cloud-based DT framework for collaborative robot management, enabling status tracking and fault detection. Similarly, Schleich et al. [7] introduced hybrid physical-virtual models for robot diagnostics and performance prediction. Despite these advancements, most DT implementations in robotics remain static or limited to visualization, with few addressing closed-loop feedback for system reconfiguration.
@@ II-A Robotic Simulation Platforms and Digital Twin Environments
Simulation platforms play a crucial role in the development and validation of DTs for robotic systems. A recent empirical study by Singh et al. [8] offers a detailed comparison of two prominent environments—Unity3D and Gazebo—through the development of a DT of an ABB IRB 1200 robotic arm. This investigation highlights both platforms’ strengths and limitations across metrics such as accuracy, latency, graphics rendering, ROS integration, cost, and scalability [8].
@@
Unity stands out for its high-fidelity real-time rendering, enabling visually immersive DTs with accurate trajectory replication. In the study, Unity produced significantly lower latency $(77.7\pm 15.7ms)$ and superior graphical realism compared to Gazebo [8]. However, Unity’s integration with ROS requires intermediate tools (e.g., ROS# or ROS–TCP–Connector), which introduces complexity and a steeper learning curve [8]. Despite this, Unity supports cross-platform deployment, including AR/VR devices, and benefits from fast prototyping workflows and robust asset libraries.
@@
On the other hand, Gazebo excels in native ROS compatibility and accurate physics simulation. The study reports latency of $108.8\pm 57.6ms$ and somewhat lower graphic fidelity in comparison to Unity [8], but emphasizes Gazebo’s no-cost, open-source nature—making it suitable for research and educational projects. Its deep sensor emulation, modular plugins, and standardized URDF-based simulations simplify DT implementation without substantial integration overhead [8].
@@
The comparative findings suggest that while Unity excels in visualization and low-latency interaction, Gazebo remains advantageous in scenarios where accurate physics simulation and ROS-native behavior are paramount. Hybrid approaches that leverage the strengths of both platforms—such as combining Gazebo’s physics engine with Unity’s rendering through frameworks like OpenUAV [9]—offer a promising direction for more robust DT implementations.
@@
Moreover, Unity’s ML-Agents toolkit [10] further extends its potential for AI-driven control and reinforcement learning, which is increasingly relevant for autonomous robotic systems. Nevertheless, Singh et al. [8] emphasize that neither Unity nor Gazebo has yet been fully exploited for enabling runtime reconfiguration of robots through DTs in dynamic smart environments. This underscores a current gap in simulation-driven reconfigurable robotics, motivating the need for new frameworks that integrate high-fidelity DTs with autonomous system adaptation. Finally, an important distinction between Gazebo and game engines like Unity, is the latter’s support for deformable mesh geometries [11], as game engines usually feature extended support for meshes, editors, importers and the like, while simpler simulation environments require extensive work on incorporating addons, and even then real-time mesh alteration and deformation is not guaranteed.
@@ II-B Trajectory Planning in Smart Robotic Systems
The first phase involves the generation of a 3D representation of the robot’s surroundings. This is commonly achieved using Simultaneous Localization and Mapping (SLAM) algorithms, which construct spatial maps based on onboard sensor data [12]. Techniques such as GMapping, Cartographer, and RTAB-Map are widely adopted for both indoor and outdoor scenarios [13]. In simulated environments, tools such as Unity3D and Gazebo are used to emulate SLAM pipelines [8], providing safe, repeatable conditions for evaluating localization accuracy and environmental complexity.
The second phase pertains to path planning, which focuses on computing a collision-free path from the robot’s current position to its goal. Classical graph-based algorithms such as A*, Dijkstra’s algorithm, and their dynamic variants like D* and D*-Lite are frequently used for global planning in known or partially known maps [14]. These algorithms are extensively supported in the Robot Operating System (ROS)  [15] through packages such as nav_core, global_planner, and move_base.
In contrast to graph-based techniques, sampling-based methods such as Rapidly-exploring Random Trees (RRT), RRT*, and Probabilistic Roadmaps (PRM) have gained prominence for motion planning in high-dimensional or continuous configuration spaces [16, 17, 18]. These methods do not require an explicit discretization of the workspace; instead, they construct a roadmap or tree by randomly sampling feasible states and connecting them through local planners. Their efficiency and scalability make them especially suitable for complex robotic systems with many degrees of freedom, such as manipulators or mobile manipulators. In the context of ROS, these planners are integrated through frameworks like OMPL (Open Motion Planning Library), which interfaces seamlessly with MoveIt! for planning in manipulation tasks [19].
The final phase—trajectory execution—translates the planned path into a sequence of velocity and position commands, constrained by the robot’s physical structure, actuator capabilities, and real-time state. A key tool widely used in this phase is the MoveIt! framework [15], which integrates motion planning, kinematics, collision avoidance, and controller management. MoveIt! interfaces with ROS and simulation tools like Gazebo to execute trajectories that account for real-world dynamics and safety constraints. Its use of sampling-based planners from OMPL enables support for a wide range of robotic configurations, from manipulators to mobile platforms.
@@ II-C Robotics Reconfiguration
Reconfiguration refers to the system’s ability to adapt its behavior in response to changes in environment, task, or internal status. While reconfiguration has been extensively studied in manufacturing  [20], agriculture  [21], and other systems, advances in complementary technologies—such as AI and communication infrastructure—now offer new perspectives to revisit and enhance this concept.
DT is such a technology, greatly complementing reconfiguration. It presents an opportunity to address the challenges of reconfigurable systems by creating a simultaneous digital environment in which the various configurations of the system can be tested, before being applied in the real-world[22]. Through DTs, it becomes feasible to simulate behavioral changes before execution, assess system-level implications, and guide reconfiguration with predictive feedback [23]. However, most DT implementations in robotics focus on monitoring or visualization tasks, rather than being actively integrated into the decision-making loop [6].
This disconnect highlights a critical research gap: current frameworks fail to fully leverage Digital Twins (DTs) for both real-time state tracking and autonomous, adaptive reconfiguration. Closing this gap is vital for robotics in smart environments, where systems must dynamically adapt to changing goals, infrastructure constraints, and multi-agent coordination demands  [24].
@@ III Digital Twin Driven Approach
The proposed DT-driven approach for robotic system reconfiguration is based on a key principle: the DT must simulate the robot’s environment as realistically as possible. While full realism isn’t always required and is potentially not in the scope of DT application, the DT should enable high-fidelity simulation with minimal developer effort  [25][26][27]. This has led to the adoption of powerful game engines for DT development. Additionally, robotics control platforms like ROS are commonly used to expose integration capabilities and companion packages for seamless connectivity between the game engine and robotic systems.
DT’s most important dimension lies in its real-time synchronization and exchange of data with its Physical counterpart, as well as the facilitation of its data driven decision making performed in the DT virtual environment  [28]. For robotic reconfiguration, the Digital Twin simulates key elements—such as mesh geometries, moving objects, and lighting conditions—to compute optimal trajectories, which are then deployed to the physical robot for execution. There is constant exchange of information across the channels connecting the physical and the digital dimensions of the application, with the one complementing the other.
@@ III-A Setting up of the DT
The proposed DT is based on the utilization of a specialized domain-specific language (DSL), which serves as a high-level configuration and modeling tool for the DT  [29]. This DSL provides a structured and expressive way to define all critical components, relationships, and behaviors of the physical system. It includes metadata and hierarchical descriptions of machines, controllers, communication interfaces, physical layouts, and process logic. The language must be expressive enough to articulate both static and dynamic aspects of the system: from the position, orientation, and dimensions of machines within a 3D spatial frame to the logical interconnections between control devices and machines. In the context of the present work, Automation ML  [30] is used to automatically configure the DT platform and create the 3D scenery in a Unity virtual environment  [31] .
@@
Based on this information, a DT generator software sets up, apart from the 3D scenery, the DT’s configuration of communication protocols and middleware layers that support real-time data exchange between physical components and their digital counterparts  [31]. This includes timing constraints and synchronization mechanisms that help maintain coherence between simulation and reality.

### v2

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. The definitive version was published in IEEE International Smart Cities Conference (ISC2) Prooceedings, DOI: 10.1109/ISC266238.2025.11293252.
@@ I Introduction
Robotic systems have achieved significant penetration in smart environments, including smart cities, smart manufacturing, and smart agriculture. Smart unmanned vehicles that deliver services inside cities, drones that monitor critical infrastructure of the city (i.e. traffic), UAVs that assist tourists, or even robotics that help pedestrians and especially persons with mobility impairments, exemplify the integration of robotics into smart city ecosystems  [tiddi2020robot]. Moreover, robots have become an integral part of modern agricultural infrastructure, automating tasks and assisting workers in their daily operations  [pal2023ai]. All these robotic systems require precise digital control, typically provided by industrial-grade controllers similar to those used in other smart infrastructures like buildings and factories. Another characteristic of smart environments is their changeability over time, especially regarding the choreographic aspect. It is common for a city to construct new buildings, roads or parks causing changes in the landscape, while this also stands true in the agricultural sector with changes in crop fields as well as relevant processing and storage infrastructures [el2021smart]. Changes in the topology of smart environments demand quick and dependable controller reconfiguration, so that they can adapt to changes and effectively plan and execute robotic system movements.
The concept of Digital Twins (DTs) has gained significant traction in the robotics domain, offering real-time synchronization between physical assets and their virtual counterparts  [mylonas2021digital]. Early applications focused on offline simulation and predictive maintenance in industrial robotics [tao2018digital], but recent work emphasizes their role in runtime monitoring, control, and autonomy. For instance, Lu et al. [lu2020digital] proposed a cloud-based DT framework for collaborative robot management, enabling status tracking and fault detection. Similarly, Schleich et al. [schleich2017shaping] introduced hybrid physical-virtual models for robot diagnostics and performance prediction. Despite these advancements, most DT implementations in robotics remain static or limited to visualization, with few addressing closed-loop feedback for system reconfiguration.
@@ II-A Robotic Simulation Platforms and Digital Twin Environments
Simulation platforms play a crucial role in the development and validation of DTs for robotic systems. A recent empirical study by Singh et al. [singh2025comparative] offers a detailed comparison of two prominent environments—Unity3D and Gazebo—through the development of a DT of an ABB IRB 1200 robotic arm. This investigation highlights both platforms’ strengths and limitations across metrics such as accuracy, latency, graphics rendering, ROS integration, cost, and scalability [singh2025comparative].
@@
Unity stands out for its high-fidelity real-time rendering, enabling visually immersive DTs with accurate trajectory replication. In the study, Unity produced significantly lower latency $(77.7\pm 15.7ms)$ and superior graphical realism compared to Gazebo [singh2025comparative]. However, Unity’s integration with ROS requires intermediate tools (e.g., ROS# or ROS–TCP–Connector), which introduces complexity and a steeper learning curve [singh2025comparative]. Despite this, Unity supports cross-platform deployment, including AR/VR devices, and benefits from fast prototyping workflows and robust asset libraries.
@@
On the other hand, Gazebo excels in native ROS compatibility and accurate physics simulation. The study reports latency of $108.8\pm 57.6ms$ and somewhat lower graphic fidelity in comparison to Unity [singh2025comparative], but emphasizes Gazebo’s no-cost, open-source nature—making it suitable for research and educational projects. Its deep sensor emulation, modular plugins, and standardized URDF-based simulations simplify DT implementation without substantial integration overhead [singh2025comparative].
@@
The comparative findings suggest that while Unity excels in visualization and low-latency interaction, Gazebo remains advantageous in scenarios where accurate physics simulation and ROS-native behavior are paramount. Hybrid approaches that leverage the strengths of both platforms—such as combining Gazebo’s physics engine with Unity’s rendering through frameworks like OpenUAV [anand2021openuav]—offer a promising direction for more robust DT implementations.
@@
Moreover, Unity’s ML-Agents toolkit [juliani2018unity] further extends its potential for AI-driven control and reinforcement learning, which is increasingly relevant for autonomous robotic systems. Nevertheless, Singh et al. [singh2025comparative] emphasize that neither Unity nor Gazebo has yet been fully exploited for enabling runtime reconfiguration of robots through DTs in dynamic smart environments. This underscores a current gap in simulation-driven reconfigurable robotics, motivating the need for new frameworks that integrate high-fidelity DTs with autonomous system adaptation. Finally, an important distinction between Gazebo and game engines like Unity, is the latter’s support for deformable mesh geometries [senthilkumar2023simulating], as game engines usually feature extended support for meshes, editors, importers and the like, while simpler simulation environments require extensive work on incorporating addons, and even then real-time mesh alteration and deformation is not guaranteed.
@@ II-B Trajectory Planning in Smart Robotic Systems
The first phase involves the generation of a 3D representation of the robot’s surroundings. This is commonly achieved using Simultaneous Localization and Mapping (SLAM) algorithms, which construct spatial maps based on onboard sensor data [cadena2016past]. Techniques such as GMapping, Cartographer, and RTAB-Map are widely adopted for both indoor and outdoor scenarios [labbe2019rtab]. In simulated environments, tools such as Unity3D and Gazebo are used to emulate SLAM pipelines [singh2025comparative], providing safe, repeatable conditions for evaluating localization accuracy and environmental complexity.
The second phase pertains to path planning, which focuses on computing a collision-free path from the robot’s current position to its goal. Classical graph-based algorithms such as A*, Dijkstra’s algorithm, and their dynamic variants like D* and D*-Lite are frequently used for global planning in known or partially known maps [koenig2002d]. These algorithms are extensively supported in the Robot Operating System (ROS)  [chitta2012moveit] through packages such as nav_core, global_planner, and move_base.
In contrast to graph-based techniques, sampling-based methods such as Rapidly-exploring Random Trees (RRT), RRT*, and Probabilistic Roadmaps (PRM) have gained prominence for motion planning in high-dimensional or continuous configuration spaces [lavalle1998rapidly, kavraki1996probabilistic, karaman2011sampling]. These methods do not require an explicit discretization of the workspace; instead, they construct a roadmap or tree by randomly sampling feasible states and connecting them through local planners. Their efficiency and scalability make them especially suitable for complex robotic systems with many degrees of freedom, such as manipulators or mobile manipulators. In the context of ROS, these planners are integrated through frameworks like OMPL (Open Motion Planning Library), which interfaces seamlessly with MoveIt! for planning in manipulation tasks [sucan2012open].
The final phase—trajectory execution—translates the planned path into a sequence of velocity and position commands, constrained by the robot’s physical structure, actuator capabilities, and real-time state. A key tool widely used in this phase is the MoveIt! framework [chitta2012moveit], which integrates motion planning, kinematics, collision avoidance, and controller management. MoveIt! interfaces with ROS and simulation tools like Gazebo to execute trajectories that account for real-world dynamics and safety constraints. Its use of sampling-based planners from OMPL enables support for a wide range of robotic configurations, from manipulators to mobile platforms.
@@ II-C Robotics Reconfiguration
Reconfiguration refers to the system’s ability to adapt its behavior in response to changes in environment, task, or internal status. While reconfiguration has been extensively studied in manufacturing  [Khan2022], agriculture  [HERNANDEZ2025110161], and other systems, advances in complementary technologies—such as AI and communication infrastructure—now offer new perspectives to revisit and enhance this concept.
DT is such a technology, greatly complementing reconfiguration. It presents an opportunity to address the challenges of reconfigurable systems by creating a simultaneous digital environment in which the various configurations of the system can be tested, before being applied in the real-world[grieves2017digital]. Through DTs, it becomes feasible to simulate behavioral changes before execution, assess system-level implications, and guide reconfiguration with predictive feedback [uhlemann2017digital]. However, most DT implementations in robotics focus on monitoring or visualization tasks, rather than being actively integrated into the decision-making loop [lu2020digital].
This disconnect highlights a critical research gap: current frameworks fail to fully leverage Digital Twins (DTs) for both real-time state tracking and autonomous, adaptive reconfiguration. Closing this gap is vital for robotics in smart environments, where systems must dynamically adapt to changing goals, infrastructure constraints, and multi-agent coordination demands  [lee2016introduction].
@@ III Digital Twin Driven Approach
The proposed DT-driven approach for robotic system reconfiguration is based on a key principle: the DT must simulate the robot’s environment as realistically as possible. While full realism isn’t always required and is potentially not in the scope of DT application, the DT should enable high-fidelity simulation with minimal developer effort  [rundel2023leveraging][clausen2022can][konig2025implementation]. This has led to the adoption of powerful game engines for DT development. Additionally, robotics control platforms like ROS are commonly used to expose integration capabilities and companion packages for seamless connectivity between the game engine and robotic systems.
DT’s most important dimension lies in its real-time synchronization and exchange of data with its Physical counterpart, as well as the facilitation of its data driven decision making performed in the DT virtual environment  [es2024methods]. For robotic reconfiguration, the Digital Twin simulates key elements—such as mesh geometries, moving objects, and lighting conditions—to compute optimal trajectories, which are then deployed to the physical robot for execution. There is constant exchange of information across the channels connecting the physical and the digital dimensions of the application, with the one complementing the other.
@@ III-A Setting up of the DT
The proposed DT is based on the utilization of a specialized domain-specific language (DSL), which serves as a high-level configuration and modeling tool for the DT  [lehner2025model]. This DSL provides a structured and expressive way to define all critical components, relationships, and behaviors of the physical system. It includes metadata and hierarchical descriptions of machines, controllers, communication interfaces, physical layouts, and process logic. The language must be expressive enough to articulate both static and dynamic aspects of the system: from the position, orientation, and dimensions of machines within a 3D spatial frame to the logical interconnections between control devices and machines. In the context of the present work, Automation ML  [faltinski2012automationml] is used to automatically configure the DT platform and create the 3D scenery in a Unity virtual environment  [alexakos2020iot] .
@@
Based on this information, a DT generator software sets up, apart from the 3D scenery, the DT’s configuration of communication protocols and middleware layers that support real-time data exchange between physical components and their digital counterparts  [alexakos2020iot]. This includes timing constraints and synchronization mechanisms that help maintain coherence between simulation and reality.